Joshua conducted an opinion poll among voters to determine if they approved the plan of biometrics there 365 who said they approved the plan. If there were 528 in the sample. What is p? and what is q?


1
Expert's answer
2022-02-02T08:43:31-0500

Solution:

"n=528"

365 who said they approved the plan.

Let p be the proportion of who approved the plan.

"p=\dfrac{365}{528}
\\ q=1-p=1-\dfrac{365}{528}=\dfrac{163}{528}"